Find solution homogeneous system of linear equations if possible

3x1+x2+x3=0

2x1+2x2-4x3=0

2x1-3x2+5x3=0


1
Expert's answer
2021-11-10T16:20:35-0500

"\Delta= \begin{vmatrix}
 3 & 1 &1\\
 2 & 2&-4\\2&-2&5
\end{vmatrix}" "=3(10-12)-1(10+8)+1(-6-4)=-34\ne0"


"\Delta _1=\begin{vmatrix}
 0 & 1&1 \\
 0 & 2&-4\\0&-3&5
\end{vmatrix}=0"


"\Delta _2=\begin{vmatrix}
 3 & 0 &1\\
 2 & 0&-4\\
2&0&5
\end{vmatrix}=0"


"\Delta _3=\begin{vmatrix}
 3 & 1&0\\
 2 & 2&0\\
2&-3&0
\end{vmatrix}=0"


The solution is


"x_1=\frac{\Delta _1}{\Delta}=0"


"x_2=\frac{\Delta _2}{\Delta}=0"


"x_3=\frac{\Delta _3}{\Delta}=0"